# M2R-Whisper: Multi-stage and Multi-scale Retrieval Augmentation for Enhancing Whisper

### v1

@@ I Introduction
The rapid advancements in deep learning have led to remarkable progress in automatic speech recognition (ASR) systems [1, 2, 3, 4, 5, 6, 7], resulting in models like OpenAI’s Whisper [5], Google’s USM [6], and META’s MMS [7] that support multilingual speech recognition and achieve near state-of-the-art performance. However, these models still encounter difficulties in low-resource settings [8, 9, 10, 11], such as recognizing diverse subdialects [12, 13, 14]. These limitations highlight the need for more flexible and adaptive methods to enhance ASR performance under such conditions.
@@
Beyond ICL, retrieval-augmented methods [24, 25, 26, 27, 28], have been widely adopted in NLP to enhance model performance without parameter updating. Techniques like $k$NN-LM [24] for language modeling and $k$NN-MT [25] for machine translation have shown significant promise. In ASR, $k$NN-CTC [29] uses Connectionist Temporal Classification (CTC) [30] pseudo labels to create speech-text key-value pairs and retrieves these labels during decoding to refine the output distribution, significantly improving performance in Chinese dialect ASR. Building on $k$NN-CTC, Zhou et al. [31] further extend this approach with a gated mechanism for zero-shot code-switching ASR. However, these methods are constrained by the quality of the pseudo labels, and $k$NN-CTC primarily addresses substitution errors, showing instability in correcting other types of recognition errors.
@@
This approach facilitates rapid domain adaptation without requiring parameter updates. Extensive experiments on AISHELL-1 [32] and KeSpeech [33] validate the effectiveness of our method, demonstrating that the integration of sentence-level and token-level datastores significantly enhances ASR performance.
@@
4. Comprehensive experiments demonstrate the effectiveness of our method on both Mandarin and Mandarin subdialect datasets.
@@ III-A Dataset
The details of each dataset we employ are shown in Table I. We conducted our experiments using two datasets: AISHELL-1 and KeSpeech.
@@ III-B Baselines
  • Whisper: We use the Large-V2 version of Whisper [5] as our baseline. Other methods incorporating Whisper are also based on the Large-V2 version.
@@ III-C Implementation details
We use the Whisper Large-V2 model as our baseline, which contains 1,550 million parameters and is trained on 630,000 hours of web-scraped speech data, demonstrating excellent performance on multilingual ASR tasks. We set retrieval neighbours $k$ to 16 for both sentence-level and token-level candidate retrieval processes. In practice, the number of prompts $n$ is capped at 10, discarding any additional retrieved samples due to Whisper’s 30-second input limit. For token-level $k$NN retrieval augmentation, the optimal value for the hyperparameter $\lambda$ is typically between 0.2 and 0.4, determined by tuning on the development set. For sentence-level candidate retrieval, we concatenate audio samples up to a maximum duration of 30 seconds to achieve the best performance All the reported results are based on greedy search decoding to ensure a fair comparison.

### v3

@@ I Introduction
The rapid advancements in deep learning have led to remarkable progress in automatic speech recognition (ASR) systems [1, 2, 3, 4, 5, 6, 7, 8], resulting in models like OpenAI’s Whisper [6], Google’s USM [7], and META’s MMS [8] that support multilingual speech recognition and achieve near state-of-the-art performance. However, these models still encounter difficulties in low-resource settings [9, 10, 11, 12, 13], such as recognizing diverse subdialects [14, 15, 16, 17]. These limitations highlight the need for more flexible and adaptive methods to enhance ASR performance under such conditions.
@@
Beyond ICL, retrieval-augmented methods [27, 28, 29, 30], have been widely adopted in NLP to enhance model performance without parameter updating. Techniques like $k$NN-LM [27] for language modeling and $k$NN-MT [28] for machine translation have shown significant promise. In ASR, $k$NN-CTC [31] uses Connectionist Temporal Classification (CTC) pseudo labels to create speech-text key-value pairs and retrieves these labels during decoding to refine the output distribution, significantly improving performance in Chinese dialect ASR. Building on $k$NN-CTC, Zhou et al. [32] further extend this approach with a gated mechanism for zero-shot code-switching ASR. However, these methods are constrained by the quality of the pseudo labels, and $k$NN-CTC primarily addresses substitution errors, showing instability in correcting other types of errors.
@@
This approach facilitates rapid domain adaptation without requiring parameter updates. Extensive experiments on AISHELL-1 [33] and KeSpeech [34] validate the effectiveness of our method, demonstrating that the integration of sentence-level and token-level datastores enhances ASR performance.
@@
4. Comprehensive experiments demonstrate the effectiveness of our method on both Mandarin and subdialect datasets.
@@ III-A Dataset
The details of each dataset we employ are shown in Table I. We conducted our experiments using two datasets:
@@ III-B Baselines
  • Whisper: We use the Large-V2 version of Whisper [6] as our baseline. Other methods incorporating Whisper are also based on the same version.
@@ III-C Implementation details
We employ Whisper Large-V2 as our baseline, which contains 1,550 million parameters and is trained on 630,000 hours of web-scraped speech data, demonstrating excellent performance on multilingual ASR tasks. We set retrieval neighbours $k$ to 16 for both sentence-level and token-level candidate retrieval processes. In practice, the number of prompts $n$ is capped at 10, discarding any additional retrieved samples due to Whisper’s 30-second input limit. For token-level $k$NN retrieval augmentation, the optimal value for the hyperparameter $\lambda$ is typically between 0.2 and 0.4, determined by tuning on the development set. For sentence-level candidate retrieval, we concatenate audio samples up to a maximum duration of 30 seconds to achieve the best performance All the reported results are based on greedy search decoding to ensure a fair comparison.